A toy manufacturer makes its own wind-up motors, which are then put into its toys While the toy manufacturing process is continuous, the motors are intermittent flow. Data on the manufacture of the motors appears below.


 Annual demand (D)-50,000 units Daily subassembly production rate = 1,000 

 Setup cost (S) = $85 per batch   Daily subassembly usage rate = 200

Carrying cost = $.20 per unit per year  


 1. To minimize cost, how large should each batch of subassemblies be? 

 2. What is the average inventory for this problem?

 3. What is the total annual inventory cost (holding plus setup) of the optimal behavior in this 3 Marks problem?
Answer #1

Answer 1:- Q* = [2DS/H(1-d/p)]0.5

Q* = [2*50000*85/H0.2(1-200/1000)]0.5

Q*=7288.7

Q* =7289 units

Answer 2:- Maximum inventory level =Q*[1-dp] =7289[1-200/1000] =5831

Average inventory =5831/2 =2915 units

Answer 3:-